The amount of coffe dispensed in a particular vending machine is normally distributed with a mean of 9 ounces and a standard deviation of 1 ounce. What percent of cups will contain more than 8.5 ounces? THe machine can hold up to10 ounces. What is the probability that your cup will runneth over ?
Answer by stanbon(75887)

The amount of coffe dispensed in a particular vending machine is normally distributed with a mean of 9 ounces and a standard deviation of 1 ounce.
-------------------
What percent of cups will contain more than 8.5 ounces?
z(8.5) = (8.5-9)/1 = -0.5
P(x > 8.5) = P(z > -0.5) = normalcdf(-0.5,100) = 0.6915
----------------------
THe machine can hold up to10 ounces. What is the probability that your cup will runneth over ?
Do the same as above for z(10)
===============
Cheers,
Stan H.
---------